Assume that 50% of people are left-handed. If we select 5 people at random, find the probability of each outcome described below, rounded to four decimal places:

a. There are some lefties (≥ 1) among the 5 people.

b. There are exactly 3 lefties in the group.

c. There are at least 4 lefties in the group.

d. There are no more than 2 lefties in the group.

e. How many lefties do you expect?

f. With what standard deviation?
Answer #1

ni 5 Pi0.50 I percentuge of left handed people, x Mumber of left handed people. fs Binomial (n=5, p=0.50) Pets) कर (org 20.11PC te 2) =1-CP (+= 3) + P (f =4) + RCA = =1-8 (+=3) + PC +3,4) =1-C0, 3125+ 0.1875] = 1- ongo orso б, уо 12.5 5 E (A) = np=5